1. A sound has an RMS pressure of 19.7 Pa, Determine the sound pressure level in dB? 


 2. A sound has an intensity of 0.38 W/m2, calculate the SPL in dB?  
Answer #1

the sound pressure level(Lp) is given by

Lp = 20* log(P / Po) = 10 * log(I / Io)

where Po = 20*10-6 Pa & Io = 10-12 W/m2